we te Mo Conceptual Questions 1. denotes answer available in Student Solutions Man 1. Is it possible to add a vector quantity to a scalar quar- tity? Explain. Ifthe co B is ze 2. Can the magnitude of a vector have a negative value 5. On a Explain. return will t top with the dimensions 1.0 m by 2.0 m. The book ends in th up at its initial position. (a) What is its displacement?cloc 3. A book is moved once around the perimeter of a table- be i (b) What is the distance traveled? learnin
Answer #1

1. No it is not possible to add a vector quantity to scalar quantity as vector have both magnitude and direction while scalar has only magnitude so as same type of quantity are added to each other i.e vector to vector and scalar to scalar.

2.The magnitude of a vector can not be negative for example let us suppose a vector quantity v=v1i+ v2j

then magnitude of v is given by=sqrt(v1^2 +v2^2)

As quantity in sqrt is always positive hence magnitude of vector can not be negative.

3.Displacement = direct distance between initial and final point

Distance = total distance travel between initial and final point

Hence in given question

Displacement =0 ( as direct distance between two point is 0)

Distance = 1+2+1+2=6m .